I am not sure of this answer. I think the answer has one real solution. Am I correct?
x^2-3x+10=0
Does it have one real solution, two real solutions or two non-real solutions.

Answer by stanbon(75887)

I think the answer has one real solution. Am I correct?
x^2-3x+10=0
-----------------
Look at the discriminant: b^2-4ac
3^2-4*1*10 = 9-40 = -31
Since the discriminant is negative there are no Real Number solutions.
====================
Cheers,
Stan H.